OpenStudy (yonatanbesrat):

a golf driving range charges a flat fee of 20 to practice and then 5.75 for a bucket of balls. write an equation that models the charges (c) in terms the number of bucket of balls (b) that you use.

OpenStudy (blazeryder):

\[c=20+5.75b\]

OpenStudy (yonatanbesrat):

expliation

OpenStudy (blazeryder):

So you know two things. The first is the person charges you $20 for the practice and then he charges you $5.75 per bucket. You want to write an equation that solves for c or total cost if you buy (b) amount of buckets. B can stand for any number excluding a negative number. You can not buy a negative number of buckets. . So the first thing you do is but C and then an equal sign. \[C = \] Now you have the starting point for your equation. Not you can build upon that. So the guys charges you $20. so you add 20 to the equation: \[C = 20\]. Now comes in the part with the variable, B or the amount of buckets. You know he charges $5.75 PER BUCKET so you can add that into your equation. \[c = 20 + 5.75\] After you do that you now need to add the variable B to the equation. You would put this with the 5.75. So the final equation would be : \[c = 20 + 5.75b\] Now if you bought 10 buckets of balls, then you would put 10 in place of b: \[c = 20 + 5.75 \times10\] Solve and you will get the total cost: \[C=$77.50\]

OpenStudy (blazeryder):

@Yonatanbesrat